# BERP: A Blind Estimator of Room Parameters for Single-Channel Noisy Speech Signals

### v6

@@ V Proposed Approach
Overview: BERP. Figure 2 shows the proposed BERP framework. We employ two separate modules: one (the unified module) processes a noisy reverberant speech signal to estimate RAPs and RGPs, corresponding to a scenario in which a primary speaker with background noise (e.g., distant speech, clapping, etc.) in the room. The other (the occupancy module) processes an occupancy reverberant speech signal to estimate instantaneous occupancy levels, assuming that the room contains multiple overlapping speakers against background noise. Before being fed into the RFE, and are transformed to time-frequency representations via featurizer.
@@ V-A Problem Formulation
Unified Module. As mentioned in Section I, we formulate simultaneous blind estimation of RAPs and RGPs as an MTL task. We can yield the following objective function:
is the trainable mapping function between the observed signal with length and the estimated room parameter , i.e., . is the corresponding loss between the estimated room parameter and the ground-truth room parameter . is the weight coefficient for each loss term. represents the shared parameters of the neural network model, rather represents the task-specific parameters for each . This objective function optimizes the parameters and to obtain Pareto optimality[45]. We use a shared encoder RFE to represent and multiple task-specific PPs as for each room parameter.
@@ V-B1 Featurizer
We use four types of featurizers to transform the observed input signals, including spectrogram, Gammatonegram, MFCC, and mel spectrogram. A Gammatonegram emphasizes the importance of low-frequency sections while a signal propagates within a room [37], whereas MFCC characterizes the shape of the spectral envelope of a reverberant signal, which is closely related to the MTF of room acoustics [92]. Instead, the mel spectrogram mimics human subjective perceptions of RACs. To clarify the potential benefits of auditory-perceptually based time-frequency representations for these blind estimation tasks, we also use spectrograms. For an input signal of length (i.e., and ), the featurizer converts it into an time-frequency representation, before passing it to the RFE for further processing. Here, is the time frame length, and is the number of frequency bins.
@@ V-B2 Room Feature Encoder
Inspired by [48, 47], we combine convolutional layers and the attention mechanism to capture local and global contexts together in time-frequency representations of input speech signals. As Wang[28] reported that acoustic information is spread throughout the entire time-frequency components of a reverberant signal, paying attention locally and globally together in a time-frequency map of a reverberant speech with a complicated waveform is well suited to learn more representative and robust features.
@@
Feedforward network. The FFN consists of two FC layers with prenorm layer normalization and a dropout layer (dropout rate: ) as the final layer. The first FC layer is an intermediate layer with 2048 dimensions, whereas the second is the output layer with 512 embedding dimensions. The architecture is detailed in Fig 2, functioning as . The activation function used after the intermediate layer is the swish function[93], defined as
@@
where is the optimal setting as 512 and . , , and are trainable weighting matrices of the query, key, and value of the attention mechanism, respectively. Specifically, the tensor that passes through the MHSA is as follows: .
@@ V-B3 Parametric Predictor
We show the mapping function of the PP from learned room features to a dedicated room parameter as:
@@ V-C1 Loss for the Unified Module
where is a ground-truth room parameter and is the estimated room parameter at each time frame . is set to 1. is the output time frame from the PP. Huber loss has dual sensitivities of minimum-variance estimation and robustness against outliers. It avoids the convergence problem of 1 loss on a small scale and contributes to preventing exploding gradients by clipping gradients exceeding [97]. By substituting Eq. (19) into Eq. (8) and setting to 1 for each , we can obtain the loss function for the unified module.
@@ V-C2 Loss for the Occupancy Module
Specifically, we use the cross-entropy (CE) to measure the difference between the probabilities of the estimated and ground-truth occupancy levels (the term in Eq. (9)), reflecting the multiclass nature of the task. The CE loss function is determined as follows:
@@ VI-A Experimental Setup
Training strategy. For the dataset for the unified module, we use 47,430 synthetic and 11,160 real reverberant audio clips. We randomly allocate 43,430 synthetic clips and 10,044 real clips for training, and 2,000 synthetic clips and 558 real clips for validation. The trained models are tested on separate sets of 2,000 synthetic clips and 558 real clips. For the occupancy module, we randomly split the total of 47,430 audio clips into 43,430 training, 2000 validation, and 2000 test datasets. The padding mask is deployed to ensure that the framework learns only the valid information across each minibatch. We used weighted random sampling on each batch to address the imbalance of synthetic and real samples in the training data. The RAdam optimizer with regularization is used [98], which possesses a learning rate warmup functionality without the risk of underfitting the regression tasks. We utilize cosine-annealing and tri-stage learning rate schedulers for the unified and the occupancy modules, respectively, to facilitate the convergence of the models toward the global optimums. We set a batch size of on an H100 GPU. Given the wide range of room volumes spanning from to , we apply logarithmic scaling to compress them, stabilizing the training process and improving the robustness of the models. Before passing through the featurizer, the amplitudes of the input signals are normalized via peak normalization.
@@
Baselines. In our comparative experiments, we evaluated the performance of our proposed method with five baseline architectures that are renowned in the domain of room parameter estimation amidst background noise: Full-CNN [33, 36, 37], CRNN [38, 39], CNN-MLP[35], TAE-CNN [29], and RE-NET [34]. We adopt the configurations of the Full-CNN and CRNN architectures based on their latest variants, as described in [33] and [38], respectively. We retrained and compared all baselines for the room parameters they are designed to estimate to ensure fair comparisons. All the baselines were sufficiently trained for convergence.
@@ VI-C1 Evaluation of RAPs and RGPs
First, we evaluated four BERP variants with different featurizers to determine which type is most effective and to clarify the contribution of auditory-inspired time-frequency representations. The comparison results on the real dataset are shown in Table I. Among four featurizers, MFCC performed the best, highlighting the potential benefits of auditory-inspired time-frequency representations. The mel spectrogram and Gammatonegram proved less effective for this task. We chose the best-performing MFCC variant to compare with the baselines. Table II presents the estimation results of BERP and the baseline methods on both real and synthetic datasets. All the models were evaluated via the same dataset described in Section IV-C1 with the data segmentation settings in Section VI-A. To assess the adaptability of the models to both simulated and real acoustical environments, evaluations were carried out separately on real and synthetic datasets. Notably, baseline methods rarely include evaluations of real recordings, making this comparison particularly significant. The results show that BERP significantly outperforms SOTA architectures in terms of the MAE both in simulated and real settings. The PCC also revealed that the estimated room parameters strongly agreed with the ground-truth ones, indicating the effectiveness of our model. The results of the MAE and PCC consistently indicate the superiority of the proposed architecture for jointly estimating RAPs and RGPs, achieving significantly better performance than the baselines in terms of the room parameters they are designed to estimate. Additionally, our approach can blindly estimate ten room parameters simultaneously with satisfactory accuracy, a capability that the baselines fail to achieve.
@@ VI-C2 Evaluation of instantaneous occupancy level
Table IV presents the evaluation results of the occupancy level estimation. As noted in Section II, to our knowledge, no recorded speech dataset is available for blind estimation of occupancy levels. Hence, we rely solely on synthetic data for this study. The creation of a real dataset is left for future work. Among the four featurizers, the MFCC performed best, underscoring the benefit of an auditory-inspired time-frequency representation. We then compared the MAE and F1 score for each occupancy level in Table V, using top-performing MFCC. The results show that the F1 score used in previous studies [25, 46] is insufficient because it considers any prediction not exact to be entirely incorrect, not accounting for how closely the estimated count aligns with the ground truth.
@@ VI-C3 Ablation Study
Separate estimation. This ablation study aims to verify our intuition that a joint estimation leads to better performance than separate estimations. Specifically, we trained separate models for each RAP and RGP and compared their performance with that of the joint estimation model. In total, we trained ten individual models, one for each RAP and RGP. The MFCC featurizer is used to train all the models. Table VI compares the performance of separate and joint models. The results confirm that the joint estimation model consistently outperforms the separate models for each room parameter, substantiating that mutual relatedness among RAPs and RGPs may benefit knowledge transfer and lead to more sufficient learning. It is also concluded that the estimation of the STI, , , , and are harder tasks that benefit significantly from the joint estimation strategy.
Auxiliary task learning. We aim to verify the hypothesis that adding more auxiliary tasks can improve overall performance, given the mutual relatedness among RAPs and RGPs (see Section I and II). In this case, rather than estimating two, four, or six room parameters as baselines do, introducing additional room parameters (adding auxiliary tasks) may lead to better performance of one or more primary tasks. In this ablation study, we trained BERP variants designed to estimate the same room parameters that the baselines estimate and compare them with the standard BERP model. The MFCC featurizer is used here to train all the models. Table VII shows the evaluation results, which confirm that incorporating more tasks leads to better performance.
In addition, we intend to highlight both the advantages of our architecture and the limitations of existing methods when jointly estimating numerous room parameters. In this case, we trained the baseline models to predict all RAPs and RGPs. The results in Table VII also indicate that the performance of the baselines significantly deteriorates on room parameters they were designed to estimate when many parameters are jointly estimated, which seems to support that their model design limits their ability to handle so many tasks.
@@
Hybrid vs End-to-end. We also explored a hybrid approach inspired by [29, 28]. Rather than predicting eight RAPs directly, we use a fewer-parameter stochastic RIR modeling and predict its parameters, which are then used to reconstruct the RIR and derive RAPs. In contrast, we estimate RGPs end-to-end. This method aims to simplify the model by reducing the number of parameters to be estimated, while improving the consistency among RAPs by deriving from the same RIR. More details are provided in the supplementary material. Table VIII shows the comparisons between the hybrid and direct architectures. The evaluation results indicate that the end-to-end architecture significantly outperforms the hybrid one. We suspect that the limitations of the hybrid approach stem from the mismatch of the RIR model with real RIRs, which likely prevents it from capturing all the complexities of real RIRs, thereby reducing the effectiveness of the hybrid approach. In addition, this hybrid approach might benefit less from auxiliary task learning, considering the parameters of the RIR model and and are mutually independent.
@@
Simulated vs Real RIR. This ablation study assesses the effectiveness of using simulated versus real RIRs for synthesizing reverberant signals, as discussed in Section II. We used an open-sourced pyroomacoustic RIR simulation algorithm [57] to generate simulated RIRs based on the same meta-information (room volume, , room dimensions, source and microphone positions, etc.) of real RIRs used to train the standard BERP model. Next, we applied the data synthesis pipeline described in Section IV-C1 to synthesize the observed reverberant signals via these simulated RIRs. We then trained two models: one using synthetic data derived from simulated RIRs and another using data derived from real RIRs. The evaluation results shown in Table IX indicate that using real RIRs to compile synthetic data offers superior performance. Notably, when generating these simulated RIRs, we did not completely replicate the detailed room geometries, material configurations, or configurations of interior objects (e.g., furniture, decorations) of real rooms where the real RIRs were measured, since accessing this meta-information is nearly impossible. To fully replicate the exact environments where the real RIRs were measured, one would have to reproduce precise room geometries and layouts and object arrangements. However, this is nearly an impractical task, given that obtaining 3D scans of various room architectures (e.g., hotels, churches, and concert halls) is often unavailable and recreating interior object settings is impractical, particularly at a large scale. Consequently, we offer a simplified baseline that still provides a meaningful insight into the realism of room simulation for training purposes.
@@
Without the parametric predictor. To dissect the contribution of the RFE and PP to the overall performance of BERP, we conducted an ablation study. We consistently used the MFCC featurizer. Table X compares the results obtained using solely the RFE with those achieved by using the full architecture equipped with the PP. The results show that the RFE contributes significantly to the performance of BERP.
@@ VII Conclusion
Discussion. This subsection recaps key points on how BERP’s model design leads to its superior performance. On the one hand, combining self-attention with convolutions enables BERP to capture acoustic features globally and locally, leading to more robust and task-generalizable shared representations over different tasks. Furthermore, by sharing representations across all tasks, the model benefits from inductive bias learning [55] and the regularization effect[45, 44, 54], leading to faster convergence, especially for harder tasks or those with fewer training examples [99, 54]. On the other hand, giving each task its own parameters mitigates gradient conflicts at the output layer, which is an issue that often arises in the single architecture baselines adopted. BERP can also predict the instantaneous occupancy level in a short time window that baselines do not offer. BERP further distinguishes itself as a sequence-to-sequence model, allowing it to handle arbitrary length inputs without explicit length alignment, an advantage over most existing approaches.
Conclusion. We propose BERP, a universal blind estimation framework that simultaneously predicts RAPs, RGPs, and instantaneous occupancy level from single-channel noisy speech. It consists of two modules: a unified module (for RAPs and RGPs) and an occupancy module (for predicting occupancy), each with distinct input signals. In the unified module, we employ a shared encoder paired with multiple task-specific predictors, whereas the occupancy module uses a classifier inspired by automatic speech recognition to estimate occupancy levels over time. These modules are combined into a single framework, bridging a critical gap, i.e., the lack of a universal solution for blind estimation of these room parameters. Evaluations on both real and synthetic data demonstrate that BERP significantly outperforms existing methods, achieving SOTA results by estimating eleven room parameters simultaneously for the first time. These findings underscore BERP’s potential across a wide range of applications in room acoustics, hearing aids, communication, and human–machine interactions.
Limitations and Future work. Except for occupancy level estimation, BERP assumes a dynamic-movement, single-source speech signal as the observed input. Future research will aim to address the blind estimation of RAPs and RPPs for multisource speech signals. Furthermore, this work assumes omnidirectional sources and receivers, which may limit its application, since [60] shows that directivity has a great impact on performance. We plan to explore this issue in future work. Another noted gap stems from the data generation for occupancy level estimation. We make some simplifications and limitations that might not completely reflect real-world settings, which would be left for future work, as well as for the collection of datasets of real recordings.

### v7

@@ V Proposed Approach
Overview: BERP. Figure 2 shows the proposed BERP approach. It comprises two networks, each trained independently for its specific task. One (the unified module) processes a noisy reverberant speech signal to estimate RAPs and RGPs, corresponding to a scenario in which a primary speaker with background noise (e.g., distant speech, clapping, etc.) within the room. The other (the occupancy module) processes a multitalker reverberant speech signal to estimate instantaneous occupancy levels, assuming that the room contains multiple overlapping speakers against background noise. Note that before being fed into the RFE, and are transformed into time-frequency representations via featurizer.
@@ V-A Problem Formulation
Unified Module. As mentioned in Section I, we formulate the blind estimation of RAPs and RGPs as a multitask learning problem. We can yield the following objective function:
is the trainable mapping function between the observed signal with length and the estimated room parameter , i.e., . is the corresponding loss between the estimated room parameter and the ground-truth room parameter . represents the shared parameters of the neural network model, whereas represents the task-specific parameters for each . is total number of room parameters. This objective function optimizes the parameters and to obtain Pareto optimality[46]. We use a shared encoder RFE to represent and multiple task-specific PPs as for each room parameter.
@@ V-B1 Featurizer
We use four types of featurizers to transform the observed input signals, including spectrogram, Gammatonegram, MFCC, and mel spectrogram. The Gammatonegram emphasizes the importance of low-frequency sections while a signal propagates within a room [38], whereas MFCC characterizes the shape of the spectral envelope of a reverberant signal, which is closely related to the MTF of room acoustics [93]. The mel spectrogram mimics human subjective perceptions of RACs. To clarify the potential benefits of auditory-perceptually based time-frequency representations for these blind estimation tasks, we also use spectrograms. For an input signal of length (i.e., and ), the featurizer converts it into an time-frequency representation, before passing it to the RFE for further processing. Here, is the time frame length, and is the number of frequency bins.
@@ V-B2 Room Feature Encoder
Inspired by [48, 49], we combine convolutional layers and the attention mechanism to capture local and global contexts together in time-frequency representations of input speech signals. As Wang[28] reported that acoustic information is spread throughout the entire time-frequency component of a reverberant signal, and paying attention locally and globally together in a time-frequency map of a reverberant speech with a complicated waveform is well suited for learning more representative and robust features.
@@
Feedforward network. The FFN consists of two FC layers with prenorm layer normalization and a dropout layer (dropout rate: ) as the final layer. The first FC layer is an intermediate layer with 2048 dimensions, whereas the second is the output layer with 512 embedding dimensions. The architecture is detailed in Fig. 2, functioning as . The activation function used after the intermediate layer is the swish function[94], defined as
@@
where is optimally set to 512, and . , , and are trainable weighting matrices of the query, key, and value of the attention mechanism, respectively. Specifically, the tensor that passes through the MHSA is as follows: .
@@ V-B3 Parametric Predictor
We show the mapping function of the PP from learned room features to a dedicated room parameter as follows:
@@ V-C1 Loss for the Unified Module
where is a ground-truth room parameter and is the estimated room parameter at each time frame . is set to 1. is the output time frame from the PP. Huber loss has dual sensitivities of minimum-variance estimation and robustness against outliers. It avoids the convergence problem of 1 loss on a small scale and contributes to preventing exploding gradients by clipping gradients exceeding [98].
@@ V-C2 Loss for the Occupancy Module
Specifically, we use cross-entropy (CE) to measure the difference between the probabilities of the estimated and ground-truth occupancy levels (the term in Eq. (9)), reflecting the multiclass nature of the task. The CE loss function is determined as follows:
@@ VI-A Experimental Setup
Training strategy. For the dataset for the unified module, we use 47,430 synthetic and 11,160 real reverberant audio clips. We randomly allocate 43,430 synthetic clips and 10,044 real clips for training, and 2,000 synthetic clips and 558 real clips for validation. The trained models are tested on separate sets of 2,000 synthetic clips and 558 real clips. For the occupancy module, we randomly split the total of 47,430 audio clips into 43,430 training, 2000 validation, and 2000 test datasets. The padding mask is deployed to ensure that the model learns only the valid information across each minibatch. We used weighted random sampling on each batch to address the imbalance of synthetic and real samples in the training data. The RAdam optimizer with regularization is used [99], which possesses a learning rate warmup functionality without the risk of underfitting the regression tasks. We utilize cosine-annealing and tri-stage learning rate schedulers for the unified and the occupancy modules, respectively, to facilitate the convergence of the models toward the global optimums. We set a batch size of on an H100 GPU. Given the wide range of room volumes spanning from to , we apply logarithmic scaling to compress them, stabilizing the training process and improving the robustness of the models. Before passing through the featurizer, the amplitudes of the input signals are normalized via peak normalization.
@@
Baselines. In our comparative experiments, we evaluated the performance of our proposed method with five baseline architectures that are renowned in the domain of room parameter estimation amidst background noise: Full-CNN [34, 37, 38], CRNN [39, 40], CNN-MLP[36], TAE-CNN [29], and RE-NET [35]. We adopt the configurations of the Full-CNN and CRNN architectures based on their latest variants, as described in [34] and [39], respectively. We retrained and compared all baselines on the room parameters they are designed to estimate to ensure fair comparisons. All the baselines were sufficiently trained for convergence.
@@ VI-C1 Evaluation of RAPs and RGPs
First, we evaluated four BERP variants with different featurizers to determine which type is most effective and to clarify the contribution of auditory-inspired time-frequency representations. The comparison results on the real dataset are shown in Table I. Among four featurizers, MFCC performed the best, highlighting the potential benefits of auditory-inspired time-frequency representations. The mel spectrogram and Gammatonegram proved less effective for this task. We chose the best-performing MFCC variant for comparison with the baselines. Table II presents the estimation results of BERP and the baseline methods on both real and synthetic datasets. All the models were evaluated via the same dataset described in Section IV-C1 with the data segmentation settings in Section VI-A. To assess the adaptability of the models to both simulated and real acoustical environments, evaluations were carried out separately on real and synthetic datasets. Notably, the baseline methods rarely include evaluations of real recordings, making this comparison particularly significant. The results show that BERP significantly outperforms SOTA architectures in terms of the MAE in both simulated and real settings. The PCC also revealed that the estimated room parameters strongly agreed with the ground truths, indicating the effectiveness of our model. The results of the MAE and PCC consistently indicate the superiority of the proposed architecture for jointly estimating RAPs and RGPs, achieving significantly better performance than the baselines in terms of the room parameters they are designed to estimate. Additionally, our approach can blindly estimate ten room parameters simultaneously with satisfactory accuracy, a capability that the baselines fail to achieve.
@@ VI-C2 Evaluation of instantaneous occupancy level
Table IV presents the evaluation results of the occupancy level estimation. As noted in Section II, to our knowledge, no recorded speech dataset is available for blind estimation of occupancy levels. Hence, we rely solely on synthetic data for this study. The creation of a real dataset is left for future work. Among the four featurizers, the MFCC performed best, underscoring the benefit of an auditory-inspired time-frequency representation. We then compared the MAEs and F1 scores for each occupancy level in Table V, using the top-performing MFCC. The results show that the F1 score used in previous studies [25, 47] is insufficient, as it treats any inexact prediction as entirely incorrect and fails to reflect how closely the estimated count aligns with the ground truth.
@@ VI-C3 Ablation Study
Separate estimation. This ablation study aims to verify our intuition that a joint estimation leads to better performance than separate estimations. Specifically, we trained separate models for each RAP and RGP and compared their performance with that of the joint estimation model. In total, we trained ten individual models, one for each RAP and RGP. The MFCC featurizer is used to train all the models. Table VI compares the performance of separate and joint models. The results confirm that the joint estimation model consistently outperforms the separate models for each room parameter, substantiating that mutual relatedness among RAPs and RGPs may benefit knowledge transfer and lead to more sufficient learning. Moreover, estimations of the STI, , , , and are harder tasks that benefit significantly from the joint estimation strategy.
Auxiliary task learning. We aim to verify the hypothesis that adding more auxiliary tasks can improve overall performance, given the mutual relatedness among RAPs and RGPs (see Sections I and II). In this case, rather than estimating two, four, or six room parameters as baselines do, introducing additional room parameters (adding auxiliary tasks) may lead to better performance of one or more primary tasks. In this ablation study, we trained BERP variants designed to estimate the same room parameters that the baselines estimate and compare them with the standard BERP model. The MFCC featurizer is used here to train all the models. Table VII shows the evaluation results, which confirm that incorporating more tasks leads to better performance.
In addition, we intend to highlight both the advantages of our architecture and the limitations of existing methods when jointly estimating numerous room parameters. In this case, we trained the baseline models to predict all RAPs and RGPs. The results in Table VII also indicate that the performance of the baselines significantly deteriorates on room parameters they were designed to estimate when many parameters are jointly estimated, seeming to support that their model design limits their ability to handle multiple tasks effectively.
@@
Hybrid vs End-to-end. We also explored a hybrid approach inspired by [28, 29]. Rather than predicting eight RAPs directly, we use a fewer-parameter stochastic RIR modeling and predict its parameters, which are then used to reconstruct the RIR and derive RAPs. In contrast, we estimate RGPs end-to-end. This method aims to simplify the model by reducing the number of parameters to be estimated, while improving the consistency among RAPs by deriving from the same RIR. Owing to page limitation, more details for the hybrid approach are provided in Sections II (architecture) and III (experimental results) of the supplementary material (see attached material below). Table VIII shows the comparisons between the hybrid and end-to-end architectures, which indicate that the end-to-end architecture significantly outperforms the hybrid architecture. We suspect that the limitations of the hybrid approach stem from the mismatch of the RIR model with real RIRs, which likely prevents it from capturing all the complexities of real RIRs, thereby reducing the effectiveness of the hybrid approach. In addition, this hybrid approach might benefit less from auxiliary task learning, considering that the parameters of the RIR model and and are mutually independent.
@@
Simulated vs Real RIR. This ablation study assesses the effectiveness of using simulated versus real RIRs for synthesizing reverberant signals, as discussed in Section II. We used an open-sourced pyroomacoustic RIR simulation algorithm [58] to generate simulated RIRs on the basis of the same meta-information (room volume, , room dimensions, source and microphone positions, etc.) of real RIRs used to train the standard BERP model. Next, we applied the data synthesis pipeline described in Section IV-C1 to synthesize the observed reverberant signals via these simulated RIRs. We then trained two models: one using synthetic data derived from simulated RIRs and another using data derived from real RIRs. The evaluation results shown in Table IX indicate that using real RIRs to compile synthetic data offers superior performance. Notably, when generating these simulated RIRs, we did not completely replicate the detailed room geometries, material configurations, or configurations of interior objects (e.g., furniture and decorations) of real rooms where the real RIRs were measured, since accessing this meta-information is nearly impossible. To fully replicate the exact environments where the real RIRs were measured, one would have to reproduce precise room geometries and layouts and object arrangements. However, this is nearly an impractical task, given that obtaining 3D scans of various room architectures (e.g., hotels, churches, and concert halls) is often unavailable and recreating interior object settings is impractical, particularly at a large scale. Consequently, we offer a simplified baseline that still provides a meaningful insight into the realism of room simulation for training purposes.
@@
Without the parametric predictor. To dissect the contributions of the RFE and PP to the overall performance of BERP, we conducted an ablation study. We consistently used the MFCC featurizer. Table X compares the results obtained using solely the RFE with those achieved by using the full architecture equipped with the PP. The results show that the RFE contributes significantly to the performance of BERP.
@@ VII Conclusion
Discussion. This subsection recaps key points related to how BERP’s model design leads to its superior performance. On the one hand, combining self-attention with convolutions enables BERP to capture acoustic features globally and locally, leading to more robust and task-generalizable shared representations over different tasks. Furthermore, by sharing representations across all tasks, the model benefits from inductive bias learning [56] and the regularization effect[46, 45, 55], leading to faster convergence, especially for harder tasks or those with fewer training examples [100, 55]. On the other hand, giving each task its own parameters mitigates gradient conflicts at the output layer, which is an issue that often arises in the single architecture baselines adopted. BERP can also predict the instantaneous occupancy level in a short time window that baselines do not offer. BERP further distinguishes itself as a sequence-to-sequence model, allowing it to handle arbitrary length inputs without explicit length alignment, an advantage over most existing approaches.
Conclusion. We propose BERP, a blind estimation approach that compromises two separately trained modules: one that jointly estimates RAPs and RGPs, and another that predicts the occupancy level, each using its own single-channel speech input. For RAP and RGP estimation, we employ a shared encoder paired with multiple task-specific predictors, whereas we use an identical encoder coupled with a classification head to estimate occupancy levels over time. They share a similar design philosophy and offer a more complete solution for the blind estimation of room acoustics. Our approach fills the current gap by providing a universal solution that addresses multiple room parameters and occupancy levels. Evaluations on both real and synthetic data demonstrate that BERP significantly outperforms existing methods, achieving SOTA results by estimating eleven room parameters simultaneously for the first time. These findings underscore BERP’s potential across a wide range of applications in room acoustics, hearing aids, communication, and human–machine interactions.
Limitations and Future work. Except for occupancy level estimation, BERP assumes a dynamically moving, single-source speech signal as the observed input. Future research will aim to address the blind estimation of RAPs and RPPs for multisource speech signals. Furthermore, this work assumes omnidirectional sources and receivers, which may limit its application, since [61] showed that directivity has a great impact on performance. We plan to explore this issue in future work. Another noted gap stems from the data generation for occupancy level estimation. We make certain simplifications and limitations that may not fully reflect real-world settings; addressing these issues, along with the collection of datasets of real recordings, is left for future work.